# A Prefrontal Cortex-inspired Architecture for Planning in Large Language Models

### v3

@@ A.6.1 Task Decomposer
Consider the following puzzle problem: Problem description:
@@
Goal: The goal is to generate a single subgoal from the current configuration, that helps in reaching the goal configuration using minimum number of moves. To generate subgoal use the goal recursion strategy. First if the smallest number isn’t at the correct position in list C, then set the subgoal of moving the smallest number to its correct position in list C.But before that, the numbers larger than the smallest number and present in the same list as the smallest number must be moved to a list other than list C. This subgoal is recursive because in order to move the next smallest number to the list other than list C, the numbers larger than the next smallest number and present in the same list as the next smallest number must be moved to a list different from the previous other list and so on.
@@ A.6.2 Actor
Consider the following puzzle problem: Problem description:
@@
A.6.3 Monitor
The $\operatorname{Monitor}$ was prompted with chain-of-thought reasoning in which each of the rules of the task were checked before determining action validity. We stored the actions deemed valid by the $\operatorname{Monitor}$ in a separate buffer, and we terminated the action proposal loop (Algorithm  1) when there were two distinct actions in this buffer, or exceeded a maximum of 10 interactions with the $\operatorname{Monitor}$. After termination of the action proposal loop, if the buffer didn’t contain two distinct actions, we used the only action in the buffer. If the buffer was empty, we used the action(s) proposed by the $\operatorname{Actor}$ at the last attempt. The following text was used as a prompt and in-context examples in the ToH task:
@@ A.6.6 Orchestrator
Consider the following puzzle problem: Problem description:
@@
A.6.7 Zero-shot prompt
@@
Consider the following puzzle problem: Problem description:
@@ A.6.8 ICL prompt
Consider the following puzzle problem: Problem description:
@@ A.6.9 CoT ICL prompt
Consider the following puzzle problem: Problem description:
@@
A.7 Computational Cost: Thinking Fast and (very) Slow
Table 7: Average per-problem computational cost ($\pm$ the standard error of the mean) on ToH with 3 disks.
Model Num. calls Num. input tokens Num. output tokens
GPT-4 ICL 1 $\pm$ 0.0 810.88 $\pm$ 0.1 190.38 $\pm$ 15.4
LLM-PFC 148.6 $\pm$ 8.2 109,090.025 $\pm$ 6,567.2 14,543.57 $\pm$ 844.6
Table 8: Average per-problem computational cost ($\pm$ the standard error of the mean) incurred by each module of LLM-PFC on ToH with 3 disks.

### v4

@@ A.7.1 Task Decomposer
Consider the following puzzle problem:
Problem description:
@@
Goal: The goal is to generate a single subgoal from the current configuration, that helps in reaching the goal configuration using minimum number of moves.
To generate subgoal use the goal recursion strategy. First if the smallest number isn’t at the correct position in list C, then set the subgoal of moving the smallest number to its correct position in list C.But before that, the numbers larger than the smallest number and present in the same list as the smallest number must be moved to a list other than list C. This subgoal is recursive because in order to move the next smallest number to the list other than list C, the numbers larger than the next smallest number and present in the same list as the next smallest number must be moved to a list different from the previous other list and so on.
@@ A.7.2 Actor
Consider the following puzzle problem:
Problem description:
@@
A.7.3 Monitor
The $\operatorname{Monitor}$ was prompted with chain-of-thought reasoning in which each of the rules of the task were checked before determining action validity. We stored the actions deemed valid by the $\operatorname{Monitor}$ in a separate buffer, and we terminated the action proposal loop (Algorithm  2) when there were two distinct actions in this buffer, or exceeded a maximum of 10 interactions with the $\operatorname{Monitor}$. After termination of the action proposal loop, if the buffer didn’t contain two distinct actions, we used the only action in the buffer. If the buffer was empty, we used the action(s) proposed by the $\operatorname{Actor}$ at the last attempt. The following text was used as a prompt and in-context examples in the ToH task:
@@ A.7.6 Orchestrator
Consider the following puzzle problem:
Problem description:
@@
A.7.7 Zero-shot prompt
@@
Consider the following puzzle problem:
Problem description:
@@ A.7.8 ICL prompt
Consider the following puzzle problem:
Problem description:
@@ A.7.9 CoT ICL prompt
Consider the following puzzle problem:
Problem description:
@@
A.8 Computational Cost: Thinking Fast and (very) Slow
Tables 16 and 17 show various cost metrics for both MAP vs. baseline models, and compares these cost metrics with performance. To address the significant computational cost of MAP, we also developed a more efficient version that cached and re-used results for redundant prompts. This was done for all modules except the $\operatorname{Actor}$ and the $\operatorname{TaskDecomposer}$. This version of the model was significantly more efficient, while retaining the same level of performance.
Table 16: Average per-problem computational cost ($\pm$ the standard error of the mean) on ToH with 3 disks. 5 runs were done for MAP.
Model Num. calls Num. input tokens Num. output tokens % solved
GPT-4 ICL 1 $\pm$ 0.0 810.88 $\pm$ 0.1 190.38 $\pm$ 15.4 46
GPT-4 CoT 1$\pm$ 0.0 1309.88 $\pm$ 0.1 422.42 $\pm$ 22.8 42
ToT 45.75 $\pm$ 0.6 26649.88 $\pm$ 454.0 7523.96 $\pm$ 322.4 6
MAP 148.6 $\pm$ 8.2 109,090.025 $\pm$ 6,567.2 14,543.57 $\pm$ 844.6 74 $\pm$ 3
MAP (efficient) 41.99 $\pm$ 3.6 47242.43 $\pm$ 3848.4 2766.34 $\pm$ 324.5 77 $\pm$ 3
Table 17: Average per-problem computational cost ($\pm$ the standard error of the mean) incurred by each module of MAP on ToH with 3 disks.